OpenStudy (anonymous):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y

OpenStudy (anonymous):

First equation, solve for y. y=2x+8. Then in the second equation, substitute in 2x+8 when you see a y.

OpenStudy (anonymous):

Let's get y by itself on the first equation. y = 8 +2x Now we substitute y in the second equation, which is 8+2x 16 + 4x = 2(8+2x) 16 + 4x = 16 + 4x Looks like it ends up cancelling both sides, so x = 0 Now we plug in 0 for x to solve for y y = 8 + 2(0) y = 8 Let's double check. 8 - 0 = 8 (Confirmed first equation) 16 - 0 = 2*8 (Confirmed second)

OpenStudy (anonymous):

16+4x=2(2x+8) then solve the equation for x

OpenStudy (anonymous):

Soo the solution is what?

OpenStudy (anonymous):

Solution is y=8 and x=0

OpenStudy (anonymous):

I don't think so. you get 16+4x=4x+16 everything cancels out so you get 0=0. I believe the solution is all real numbers

OpenStudy (anonymous):

16+4x-4x=4x+4x+16 16=16 or 16-16+4x-4x=4x-4x+16-16 0=0

OpenStudy (anonymous):

My options are the solution is (24,0), the solution is (1,10), there is no solution, and There are an infinite number of solutions

OpenStudy (anonymous):

My bad, it's infinite number of solutions as DSS said.

OpenStudy (anonymous):

Thanks G-unit and DSS!!!